# Self-Calibrating Gaussian Splatting for Large Field of View Reconstruction

### v1

@@ 4 Optimization of Camera Parameters
In this section, our method can accurately recover the camera parameters of a perspective pinhole camera model. We use the NeRF-Synthetic dataset [48], which includes known ground truth camera poses. The dataset contains 100 viewpoints for training and 200 additional viewpoints for computing test error metrics. We first add noise to perturb the rotation angles of camera extrinsics, the positions of the camera centers, and the focal lengths. These noisy cameras are used to train both the baselines and our methods. We compare our methods with CamP [52] implemented on ZipNeRF [7], a state-of-the-art method for joint optimization of 3D scene and camera extrinsics and intrinsics. In addition to CamP, we also report the performance of vanilla Gaussian Splatting without pose optimization.
The models are evaluated on two criteria, following the protocol in CamP [52]. First, we measure the accuracy of the estimated camera pose in the training views after accounting for a global rigid transformation. Second, we measure the quality of rendered images at the held-out camera poses after a test-time optimization to adjust the held-out pose for a potential global offset. Tab. 5 shows that our method can outperform both vanilla 3DGS and CamP in both image and camera metrics by a large margin.
@@
We further test our model by perturbing the camera poses with different levels of noise. Specifically, we add Gaussian noise with a standard deviation ranging from to to the camera poses. Adding larger noise to camera poses is very challenging. Fig. 10 and Fig. 11 show that CamP’s performance signifantly drops since it is prone to falling into local minima when the initial camera poses have significant errors, while our method shows much slower degradation in novel-view synthesis performance.
(a) Lego | (b) Drums | |||
(c) Material | (d) Chair |
@@ 5 Computational Efficiency
To verify the hypothesis that our hybrid method allows us to find a better combination of expressivity and speed, we perform an ablation study on the FisheyeNeRF dataset [34]. Specifically, we study the grid resolution for , which determines the number of evaluations needed for the most expensive part of the pipeline – the invertible ResNet. Table 6 reports the PSNR and training time for three different grid resolutions, Fisheye-GS [46], and vanilla 3DGS [37]. We can see that higher resolution leads to better performance but longer training time. Reducing the grid resolution further does not lead to shorter computation time, since other operations, such as the gradient computation for other camera parameters, start to dominate. Fortunately, all hybrid solutions significantly outperform vanilla 3DGS and Fisheye-GS. The lowest resolution we tried introduces only minutes of training time overhead on top of 3DGS, in exchange for a dB boost in PSNR.
@@
A higher resolution control grid results in a smoother distortion field representation. To better illustrate the effect of control grid resolution, we visualize two types of resolutions (i.e., 265 149 and 66 37). While there are no significant differences in the central region, the distortion at the edges is better recovered with a higher resolution grid. Since the distortion field becomes smoother, a high-resolution control grid produces more accurate distorted lines, as shown in the red and blue boxes of Fig. 9.
(a) Original | (b) Cropped |
@@
(a) Living | (b) Kitchen | (c) Room |
6 Real-world Walk-around Scenes
We use a Canon 5D Mk III DLSR camera with a Canon 8mm-15mm Fisheye lens zoomed out to 8mm with a 180 FOV to capture a complex indoor office scene, where we place models and spheres on a table. Images are taken close to the table to capture the details of the various models and spheres. We also use Meike 3.5mm f/2.8 Ultra Wide-Angle Circular Fisheye Lens capture the same office. However, in practice, we did not fully utilize the entire 220 field of view, primarily because the ultra-wide angle often includes the tripod at the bottom (handholding is impractical, as the human body would also appear in the lower region). Moreover, when the raw 220 images are fed into COLMAP [58], the SfM pipeline fails because this camera captures scenes from the back, and the projection models predefined in COLMAP [58] do not support back-facing projections. Consequently, we crop the raw images slightly to reduce the FOV to 180 and partially exclude the tripod’s influence like Fig. 12. Qualitative results are shown in Fig. 13. Our method effectively recovers details in the central region while producing accurate lens distortion for background elements, such as painting frames and lines on the white wall.
@@ 6 Real-world Walk-around Scenes
We also mount two fisheye cameras on a rig configured such that the cameras are perpendicular to each other. This camera rig is used to capture a backyard scene by walking clockwise and counterclockwise twice to record videos. The benefit of using this rig is that the relative pose between the two fisheye cameras is fixed, simplifying the SfM [58] pipeline for estimating accurate poses. As shown in Fig. 14, our method converges to an accurate calibration, ensuring that lines on the house’s surface and the ladder leaning on the tree are straight when rendered in perspective views.
7 Additional Ablation Studies
Ablation of Cubemap
As illustrated in Fig. 2 of the main paper, we apply cubemap to deal with the limitation of perspective projection. Even with our hybrid distortion field, we can only take advantage of center part of raw fisheye images as shown in the bottom row of Fig. 15. In contrast, rendering cubemap allows us to expand to larger FOV and thus compute the photometric loss with reference raw images for optimization like Fig. 5 (d) in the main paper. Because of the rendering FOV of single planar projection and cubemap, the final reconstruction quality is compromised as shown in Tab. 3 of the main paper.
Besides, the boundary of the final distorted rendering is irregular, largely because the distortion information at the boundary heavily relies on the extra regions that are not covered by single planar projections.
Ablation of Hybrid Field Optimization.
In practice, the estimated distortion from the SfM [58] pipeline can be used as an initialization for our hybrid field, which makes training stable and achieves faster convergence. However, these parameters may be inaccurate when solving from highly distorted images. We verify the necessity of optimizing our hybrid field during reconstruction both quantitatively and qualitatively. If we assume the distortion parameters are accurate and expressive enough to model lens distortion, we can freeze the hybrid field after initialization from SfM. While the hybrid field can produce a roughly correct distortion field, it leads to blurry reconstructions, as shown in Fig. 16. This degradation becomes more pronounced in scenes with many straight lines, such as the jalousie windows in chairs scene. The photometric evaluation in Tab. 7 shows our final hybrid field can outperform estimated distortion parameters by a large margin. Optimizing our hybrid field is crucial for achieving more accurate reconstruction and distortion modeling.
Method | Chairs | Cube | Flowers | Globe | Heart | Rock | ||||||||||||
COLMAP+GS | 0.583 | 18.28 | 0.290 | 0.637 | 21.64 | 0.296 | 0.443 | 18.09 | 0.379 | 0.580 | 19.63 | 0.327 | 0.660 | 20.87 | 0.282 | 0.511 | 20.24 | 0.280 |
Before Opt | ![]() |
After Opt | ![]() |
GT | ![]() |
(a) Chairs | (b) Cube | (c) Flowers |
Ablation Effective Regularization of iResNet.

### v2

@@ 2 Optimization of Camera Parameters
In this section, we first derive the gradients for all camera parameters during training in Sec. 2.1. We then demonstrate the effectiveness of the joint optimization of distortion alongside extrinsic and intrinsic parameters in Sec. 2.2. Finally, we evaluate our camera optimization on synthetic NeRF [50] scenes with significant noise in Sec. 2.3.
2.1 Derivation of Gradient
We first derive how to compute the gradient for a pinhole camera and then extend the derivation to account for distortion.
As defined above, the gradient of camera parameters can be written as:
For the color term, we first define the input of the spherical harmonics stored by each Gaussian as the view direction. The 3D location of a Gaussian in world coordinates is , and the camera center is :
Then, the derivative with respect to the camera center is:
For now, we only compute the derivative with respect to the camera center. However, the gradient can be easily propagated back since the camera center corresponds to the translation vector of the inverse of the world-to-camera matrix.
Next, we derive the gradient from the projected 2D Gaussian position . For simplicity, we integrate the intrinsic and extrinsic parameters into a projection matrix P since we do not yet consider distortion. The projection from a 3D Gaussian in world coordinates is defined by:
We apply a projective transformation:
where
Then, we project from NDC to screen space to obtain the 2D location :
The second term can be represented as:
We compute the gradients for each in the projection matrix P:
and
The gradient flow back to intrinsic and extrinsic parameters can be easily computed since .
Finally, we compute the last term. The 2D covariance depends only on the view matrix, so instead of using P, we use only the view matrix (i.e., the world-to-camera matrix) V, and the transformed 3D location in camera coordinates is :
We compute the 2D covariance as follows, given known focal lengths and :
W | ||||
T | ||||
Conic |
Since we only extract the upper three elements of , we do not compute gradients for , , and .
To account for distortion, we decompose the projection matrix into two separate operations. Following the definitions above, we apply an invertible ResNet in between:
For the projection of the Gaussian covariance , we compute a new Jacobian matrix using the distorted :
Both processes are differentiable, allowing us to compute the intermediate Jacobian for to update the invertible ResNet.
Chairs | ![]() |
Rock | ![]() |
Chairs | ![]() |
Globe | ![]() |
Flowers | ![]() |
Flowers | ![]() |
(a) w/o Optimize Cameras | (b) Optimize Cameras | (c) GT |
(a) Lego | (b) Drums | |||
(c) Material | (d) Chair |
2.2 Joint Distortion and Pose Optimization
Our approach supports efficient optimization of camera parameters, either independently (with perspective images) or in combination, as shown in Fig. 8, with distortion modeling. This ensures that our pipeline remains robust even when both distortion and camera parameters are inaccurate.
To evaluate our model’s ability to jointly optimize lens distortion and other camera parameters in real-world scenes, we introduce a setting where both camera extrinsics and intrinsics are perturbed in the FisheyeNeRF dataset. Specifically, we add Gaussian noise with a standard deviation of to each image’s camera extrinsic and intrinsic parameters. Despite this additional noise, our method successfully recovers the lens distortions while generating high-quality novel-view synthesis renderings (Fig. 10).
One key observation is that COLMAP [59] can robustly estimate camera extrinsics but struggles with intrinsic parameters and distortion. When we enable extrinsic optimization during reconstruction, the camera poses are refined only slightly, indicating that the initial poses are already quite reliable. Regarding lens distortion, as demonstrated in Fig. 12 and discussed in Hybrid Field of Sec.4.4 andSec. 3, COLMAP’s distortion estimation lacks accuracy, highlighting the necessity of our hybrid distortion field for improved expressiveness and precision.
We hypothesize that this limitation stems from the structure-from-motion (SfM) pipeline [59] in COLMAP. COLMAP primarily utilizes the central region of raw images (corresponding to a small FOV), where conventional distortion models perform well. Consequently, lens distortion has minimal influence on extrinsic estimation, as COLMAP can still rely on the image center to solve linear equations. However, when attempting to leverage the full FOV of raw images for reconstruction, the limitations of a fixed distortion model and a single-plane projection become pronounced.
To validate the self-calibration capability of our pipeline, we manually introduce noise into the extrinsics produced by COLMAP and jointly optimize extrinsics, intrinsics, and distortion. As shown in Fig. 10, our approach effectively refines all camera parameters and distortion simultaneously. Without extrinsic and intrinsic optimization, the hybrid field can only predict coarse distortion, while misaligned poses contribute to the blurry reconstruction seen in Fig. 10 (a). Even with significant perturbations in poses and intrinsics, our method robustly recovers accurate camera parameters and the distortion field after training, as illustrated in Fig. 10 (b).
@@ 2.2 Joint Distortion and Pose Optimization
Scenes | Metric | 0 | 0.1 | 0.15 | 0.2 | 0.25 |
Chair | SSIM | 0.987 | 0.988 | 0.988 | 0.987 | 0.980 |
PSNR | 35.81 | 36.10 | 35.99 | 35.83 | 34.35 | |
LPIPS | 0.012 | 0.012 | 0.012 | 0.012 | 0.019 | |
Lego | SSIM | 0.983 | 0.974 | 0.972 | 0.969 | 0.965 |
PSNR | 35.77 | 34.78 | 34.29 | 33.70 | 33.15 | |
LPIPS | 0.015 | 0.021 | 0.023 | 0.025 | 0.029 | |
Drums | SSIM | 0.955 | 0.954 | 0.953 | 0.953 | 0.946 |
PSNR | 26.17 | 26.15 | 26.04 | 26.01 | 25.30 | |
LPIPS | 0.037 | 0.038 | 0.039 | 0.040 | 0.045 | |
Materials | SSIM | 0.960 | 0.960 | 0.951 | 0.942 | 0.843 |
PSNR | 29.99 | 29.93 | 28.91 | 27.95 | 15.17 | |
LPIPS | 0.034 | 0.036 | 0.044 | 0.052 | 0.158 | |
Mic | SSIM | 0.991 | 0.989 | 0.987 | 0.974 | 0.923 |
PSNR | 35.34 | 34.58 | 33.65 | 30.07 | 18.78 | |
LPIPS | 0.006 | 0.008 | 0.010 | 0.019 | 0.087 | |
Ship | SSIM | 0.907 | 0.873 | 0.776 | 0.719 | 0.700 |
PSNR | 30.91 | 28.66 | 20.96 | 16.80 | 15.10 | |
LPIPS | 0.106 | 0.126 | 0.203 | 0.262 | 0.294 | |
Ficus | SSIM | 0.987 | 0.987 | 0.984 | 0.955 | 0.859 |
PSNR | 34.85 | 34.84 | 33.99 | 28.08 | 18.54 | |
LPIPS | 0.012 | 0.012 | 0.014 | 0.039 | 0.125 | |
Hotdog | SSIM | 0.985 | 0.985 | 0.985 | 0.983 | 0.982 |
PSNR | 37.67 | 37.65 | 37.63 | 37.05 | 36.53 | |
LPIPS | 0.020 | 0.020 | 0.020 | 0.022 | 0.025 |
Method | Chairs | Cube | Flowers | Globe | Heart | Rock | ||||||||||||
Ours (Freeze Init) | 0.583 | 18.28 | 0.290 | 0.637 | 21.64 | 0.296 | 0.443 | 18.09 | 0.379 | 0.580 | 19.63 | 0.327 | 0.660 | 20.87 | 0.282 | 0.511 | 20.24 | 0.280 |
Chairs | ![]() |
Cube | ![]() |
Flowers | ![]() |
(a) Freeze Intialization | (b) Ours | (c) GT |
2.3 Pure Camera Optimization
We use the NeRF-Synthetic dataset [50], which includes known ground truth camera poses. The dataset contains 100 viewpoints for training and 200 additional viewpoints for computing test error metrics. We first add noise to perturb the rotation angles of camera extrinsics, the positions of the camera centers, and the focal lengths. These noisy cameras are used to train both the baselines and our method. We compare our method with CamP [54], implemented on ZipNeRF [7], a state-of-the-art method for joint optimization of 3D scenes and camera extrinsics and intrinsics. In addition to CamP, we also report the performance of vanilla Gaussian Splatting without pose optimization.
The models are evaluated on two criteria, following the protocol in CamP [54]. First, we measure the accuracy of the estimated camera poses in the training views after accounting for a global rigid transformation. Second, we measure the quality of rendered images at the held-out camera poses after a test-time optimization to adjust for a potential global offset. Tab. 5 shows that our method outperforms both vanilla 3DGS and CamP in both image and camera metrics by a large margin.
We further evaluate our model by perturbing the camera poses with varying levels of noise. Specifically, we add Gaussian noise with a standard deviation ranging from 0 to 0.3 to the camera poses. For reference, we report the increased noise levels for each scene in Tab. 6. Handling larger noise in camera poses presents a significant challenge. As shown in Fig. 9 and Fig. 11, CamP’s performance drops significantly due to its tendency to fall into local minima when the initial camera poses contain substantial errors. In contrast, our method exhibits much slower degradation in novel-view synthesis performance.
We also include the video “pose_opt.mp4” to visualize the optimization process.
3 Distortion Estimation from COLMAP
In practice, the distortion estimated from the SfM [59] pipeline can be used as an initialization for our hybrid field, stabilizing training and accelerating convergence. However, these parameters are inaccurate when derived from highly distorted images. We verify the necessity of optimizing our hybrid field during reconstruction both quantitatively and qualitatively.
To assess this, we fit the COLMAP [59] distortion parameters within our hybrid field and freeze the network. As shown in Tab. 7, the distortion initialization from COLMAP improves upon Vanilla 3DGS [38] but remains far from accurate in modeling distortion compared to our approach. While the hybrid field produces a roughly correct distortion field, it results in blurry reconstructions, as shown in Fig. 12. This degradation is particularly noticeable in scenes with many straight lines, such as the jalousie windows in the chairs scene. The photometric evaluation in Tab. 7 demonstrates that our fully optimized hybrid field significantly outperforms the estimated distortion parameters. These results highlight the importance of optimizing our hybrid field for achieving more accurate reconstruction and distortion modeling.
@@ 4 Computational Efficiency
To verify the hypothesis that our hybrid method achieves a better balance between expressiveness and efficiency, we perform an ablation study on the FisheyeNeRF dataset [36]. Specifically, we analyze the grid resolution of , which determines the number of evaluations required for the most computationally expensive part of the pipeline—the invertible ResNet. Tab. 8 reports the PSNR and training time for three different grid resolutions, as well as for Fisheye-GS [47] and vanilla 3DGS [38]. We observe that increasing the resolution of leads to better performance but longer training time. Further reducing the grid resolution does not significantly shorten computation time, as other operations, such as gradient computation for camera parameters, begin to dominate. Fortunately, all hybrid solutions significantly outperform vanilla 3DGS and Fisheye-GS. The lowest-resolution setting we tested introduces only a -minute training time overhead compared to 3DGS, in exchange for a dB boost in PSNR.
(a) Low Resolution | (b) High Resolution |
@@
A higher-resolution control grid results in a smoother distortion field representation but slower training. To better illustrate the effect of control grid resolution, we visualize two types of resolutions (i.e., 265 149 and 66 37). While there are no significant differences in the central region, the distortion at the edges is better recovered with a higher-resolution grid. Since the distortion field becomes smoother, a high-resolution control grid produces more accurate distorted lines, as shown in the red and blue boxes of Fig. 13.
5 Extra Experiments
Method | Garden | Studio | ||||
SSIM | PSNR | LPIPS | SSIM | PSNR | LPIPS | |
Fisheye-GS [47] | 0.530 | 14.94 | 0.542 | 0.536 | 12.24 | 0.549 |
Ours | 0.882 | 27.85 | 0.144 | 0.965 | 33.86 | 0.044 |
Method | Test View | Num | SSIM | PSNR | LPIPS |
Fisheye-GS [47] | Fisheye | 100 | 0.630 | 15.94 | 0.463 |
Ours | 100 | 0.886 | 30.72 | 0.146 |
5.1 Quantitative Comparisons with Fisheye-GS
In addition to Fig. 5 in the main paper, we also provide a quantitative evaluation of our method compared with the baseline Fisheye-GS [47] in Tab. 9 and Tab. 10. The performance degradation observed in the baseline method is primarily due to the limitations of the conventional camera distortion model used during reconstruction, which struggles at the edges of large FOVs. As a result, there is no geometric consistency in the peripheral regions to produce uniform gradients for optimizing the Gaussians. When rendering novel views, these Gaussians appear as large floaters that occlude the camera view. We further visualize this phenomenon in the failure case video reconstructed by Fisheye-GS [47]. The center area is revealed as we gradually decrease the scale of the Gaussians. Please refer to our supplementary video “fisheye-gs_failure.mp4” for a comparison with the parametric distortion model.
(a) Living | (b) Kitchen | (c) Room |
5.2 Qualitative Results of Cubemap Ablation
As illustrated in Fig. 2 of the main paper, we apply a cubemap to overcome the limitations of perspective projection. Even with our hybrid distortion field, we can only utilize the central region of raw fisheye images, as shown in the bottom row of Fig. 14. In contrast, rendering a cubemap enables us to achieve a larger FOV, allowing us to compute the photometric loss with reference raw images for optimization, as demonstrated in Fig. 7 of the main paper. We also provide a quantitative evaluation in Tab. 4 of the main paper, showing that the final reconstruction quality is compromised when using a single planar projection.
Additionally, the boundary of the final distorted rendering is irregular, primarily because the distortion information at the boundary heavily relies on extra regions that are not covered by single planar projections.
5.3 Qualitative Comparisons with Different Numbers of Input Views
We report quantitative results in Tab. 2 of the main paper, where our method outperforms the baseline even with significantly fewer input views, down to 10-15%.
To illustrate the impact of varying input view counts, we visualize the rendering quality as the number of input views decreases. Even with as few as 25 input views, our method still achieves reasonable performance, as shown in Fig. 15. This demonstrates our method’s ability to effectively utilize large FOV cameras and achieve comprehensive scene coverage during reconstruction.
(a) 100 Views | (b) 50 Views | (c) 25 Views | (d) 10 Views |
5.4 Comparisons with Regular FOV Cameras
Synthetic Mitsuba Scene Captures.
To carefully control the experimental settings, we customized a camera module in the Mitsuba ray tracer [35] using camera parameters derived from DSLR lenses, as profiled in the open-source Lensfun [1]. We also utilize 3D assets, including geometry, materials, and lighting, from [9] to produce renderings. Our synthetic dataset contains three large indoor scenes and four object-centric scenes. All indoor scenes follow a Sigma 180 circular fisheye camera. Two of the object scenes are rendered with a 120 fisheye lens, while the others are rendered with classic radial distortion.
To capture detailed perspectives of the scene, cameras are placed close to the objects at varying distances. For Room2, since objects are uniformly distributed in the space, we placed a set of camera centers along a Hilbert curve, then oriented each fixed camera center to cover the surroundings. The number of images captured at each point is reduced for 180 images compared to those with a 90 FOV, as shown in the number of captures in Tab. 11.
Method | Num | Mitsuba Kitchen | Mitsuba Room1 | Mitsuba Room2 | ||||||
SSIM | PSNR | LPIPS | SSIM | PSNR | LPIPS | SSIM | PSNR | LPIPS | ||
3DGS [38] | 200 | 0.470 | 11.09 | 0.435 | 0.595 | 15.27 | 0.406 | 0.897 | 30.88 | 0.155 |
Ours | 10 | 0.644 | 23.03 | 0.346 | 0.556 | 22.26 | 0.377 | 0.644 | 23.03 | 0.346 |
25 | 0.689 | 24.75 | 0.303 | 0.613 | 24.21 | 0.318 | 0.825 | 24.81 | 0.254 | |
50 | 0.708 | 25.44 | 0.285 | 0.640 | 25.07 | 0.301 | 0.856 | 28.26 | 0.214 | |
100 | 0.794 | 27.56 | 0.272 | 0.686 | 26.27 | 0.314 | 0.920 | 33.21 | 0.107 | |
Fisheye-GS [47] | 100 | 0.601 | 14.30 | 0.485 | 0.549 | 15.54 | 0.548 | 0.739 | 17.97 | 0.355 |
Ours | 10 | 0.801 | 26.04 | 0.202 | 0.703 | 23.69 | 0.232 | 0.806 | 19.06 | 0.256 |
25 | 0.855 | 28.62 | 0.155 | 0.783 | 26.71 | 0.180 | 0.853 | 24.28 | 0.167 | |
50 | 0.870 | 29.55 | 0.151 | 0.801 | 27.57 | 0.172 | 0.883 | 27.29 | 0.133 | |
100 | 0.886 | 30.72 | 0.146 | 0.842 | 28.46 | 0.180 | 0.929 | 31.16 | 0.095 |
Room1 | ![]() |
Kitchen | ![]() |
Room2 | ![]() |
(a) Small FOV | (b) Ours | (b) GT |
@@ 5.4 Comparisons with Regular FOV Cameras
Qualitative Results.
To verify the accuracy of self-calibration, we generate a set of hold-out cameras that share the same distribution as the training set. For each validation camera, we render paired perspective and 180 fisheye images. Unlike real-world datasets such as Garden and Studio, our synthetic dataset allows direct comparison with ground-truth perspective views. As shown in Fig. 16, we render an additional 20 for the hold-out cameras to highlight the difference in coverage between our method and conventional capture approaches.
It is worth noting that in the comparison between our method and 3DGS [38], we evaluate in perspective views since 3DGS [38] does not support fisheye rendering, whereas our method can generate perspective views after training. In contrast, we directly compare our method with Fisheye-GS [47], as both methods natively support fisheye rendering.
5.5 More Real-world Scenes
Scene Captures.
We use a Canon 5D Mk III DSLR camera with a Canon 8mm-15mm fisheye lens, zoomed out to 8mm with a 180 FOV, to capture a complex indoor office scene, where we place models and spheres on a table. Images are taken close to the table to capture the details of the various models and spheres. We also use a Meike 3.5mm f/2.8 ultra-wide-angle circular fisheye lens to capture the same office.
@@ 5.5 More Real-world Scenes
Additionally, we mount two fisheye cameras on a rig configured such that the cameras are perpendicular to each other. This camera rig is used to capture a backyard scene by walking clockwise and counterclockwise twice to record videos. The benefit of using this rig is that the relative pose between the two fisheye cameras is fixed, simplifying the SfM [59] pipeline for estimating accurate poses.
Qualitative Results.
Qualitative results are shown in Fig. 17. Our method effectively recovers details in the central region while accurately modeling lens distortion for background elements, such as painting frames and lines on the white wall.
As shown in Fig. 18, our method converges to an accurate calibration, ensuring that lines on the house’s surface and the ladder leaning against the tree remain straight when rendered in perspective views.
5.6 Adaptability to Different Lens Distortions
Method | PSNR | SSIM | LPIPS |
Vanilla-GS [38] | 25.37 | 0.923 | 0.121 |
Ours | 28.00 | 0.932 | 0.093 |
As mentioned in the last part of Sec. 4.2 of the main paper and also shown in Fig. 6, we introduced synthetic distortions, including both radial and tangential components, to images from the LLFF dataset [49].
We also apply moderate radial distortion to our synthetic dataset and reconstruct several object scenes. After training, we can render undistorted images. The perspective rendering increases the FOV while maintaining the same camera extrinsics. As shown in Fig. 19, distorted edge lines, such as those on the Lego and Car objects, are correctly recovered into straight lines, demonstrating the capability of our hybrid field to model radial distortion effectively. We also report quantitative evaluations in Tab. 12. Note that the radial distortion is relatively subtle, so the improvement is not as pronounced compared to other types of lenses.
(a) Lego | (b) Car | ||
(c) Rover | (d) Spaceship |
5.7 Comparison of Regular and Invertible ResNet